The HIRES web pages have been reformatted and reorganized to conform to the Keck instrument page standards. Feedback is welcome.
MAGIQ Guide Camera Installed (March 2012)
The original Photometrics guide camera has been replaced by a new SciMeasure Analytical Systems Little Joe CCD
camera. The detector is a 1024 x 1024 pixel, thinned, backside-illuminated CCD with a quantum effeciency
of nearly 85% in V-band. The new pixel scale is 0.085" / pixel (compared with 0.15" / pixel for the Photometrics
camera) yielding a maximum field of view of approximately 87" X 87". The original guide camera filter wheels
and filters (photometric and neutral density) remain in place as well as the Canon 200 mm f/1.8 lens which serves
as a focal reducer. More information can be found here.
Server Upgrade (April 2011)
The HIRES server (lehoula) has been upgraded from a 280R to a V240 platform. The disk
cabinets have also been replaced by a Triton 8SA storage system.
VNC Operation (June 2010)
Beginning June 2010, we ask that all observers start up
their HIRES control software within VNC desktops. Virtual
Network Computing (VNC) is a technology that allows windows to be
shared with observers at other computers and has been used
successfully for years at Keck to provide mainland observers with
access to Keck instrument controls. By expanding the use of VNC to
all observers, we will thus have a single startup procedure for
observers to follow regardless of whether or not they will operate
with mainland observers. Operating under VNC also makes it easier
for support staff to assist observers with troubleshooting and
allows our observing assistants to monitor the status of
observations during the night. Your support astronomer will be
happy to assist you with starting your HIRES software under VNC.
Bias Level Variation (February 2010)
Transient, low-level (0.5-3 DN) bias jumps have been noted in all 3 chips of the
array. The effect is currently under investigation and may impact a significant
fraction of images taken (~10%).
Shutter replaced (October 2009)
The shutter was replaced after indications that it was failing to fully open.
The new shutter is operating nominally.
Makee version 5.2.4 now available (October 31, 2008)
The latest version of Makee has been released. This version allows KOA tables
for single CCD HIRES data. It also includes expanded wavelength calibration,
especially in the red for the 3 CCD system. Makee 5.2.4 is available from the
command line for HIRES observers.
Exposure time calculator now available (September 23, 2008)
Jason Prochaska of UCO/Lick Observatory has created an exposure time calculator
for HIRES. We provide this tool with the caveat that we have not as of this
writing, had a chance to thoroughly test its output in all regimes. The ETC
is linked from the HIRES Pre-Observing page
Faster access to data through KOA (September 2, 2008)
The Keck Observatory Archive (KOA) is now transferring HIRES data
electronically to MSC for archiving. This has significantly sped up the
data ingestion process, and observers should have access to their data
through KOA well within 12 hours of observing.
New default binning (April 30 2008)
We are unaware of any HIRES observers that do not bin in the spatial direction.
The majority use 2X1 binning so that is now the default.
New default spectrograph focus (April 16 2008)
HIRES may be very slowly relaxing towards pre-earthquake focus values. The new,
default value of cafraw is -35,000 encoder counts.
Shutter repaired (February 1, 2008)
The shutter problem noted below has been fixed. Troubleshooting was complicated
by the existence of multiple, subtle problems. The position switch on the
shutter itself was intermittent and a new shutter has been installed. An IC
within the shutter controller itself needed to be replaced.
New Th/Ar comparison lamp installed April 2, 2007)
There are two Th/Ar lamps installed in HIRES. Historically, they have exhibited
excessively strong argon lines at long wavelengths. Some time ago, in an
effort to offer some alternative to observers we replaced Th/Ar number 1 with
a lamp that had some neon gas. This did dim the argon lines down but at the
expense of very strong red neon lines. We have now replaced this lamp with one
with less neon. The neon lines are now weak and although the argon lines are
weaker than in lamp number 2 they are still quite strong. Lamp number 1 does
not illuminate the longer deckers as well as lamp number 2 so observers are
encouraged to experiment and pick which lamp they want to use, depending on
their wavelength region and chosen decker length.
Shutter appears to occasionally fail to open (Aug 21 2006)
There have been increasing reports recently of the HIRES shutter
appearing as if it has failed to open.
Note; by shutter we do not mean either the hatch or
the dark slide. It may not be obvious from examination of XHIRES but there
is an actual shutter inside HIRES. The colour coded exposure box on the HIRES
dashboard remains black during a dark exposure indicating a closed shutter.
From what we can tell, the shutter is actually opening when commanded to, but
we continue to investigate.
HIRES dewar thermal cycled to replace ion pump (May 4 2006)
Over the past week, work was done to thermally cycle the HIRES
dewar, replace a failing ion pump, oxygen soak the detector and cool it back
down. The procedure appears to have been successful, and the QE recovered to
the best of our ability to determine so using flat lamps. We have redetermined
various parameters of the CCD and the user notes for the detector upgrade has
been updated. This document, available, on the HIRES pre-observing, and
observing web pages should be consulted. Most parameters only changed
slightly or not at all. Linearity remains to be remeasured.
Wavelength instability during telescope slews? (Apr 17 2006)
Some evidence exists that wavelength settings may shift by about a
pixel during telescope slews. Further testing is needed but preliminary
indications are that this may or may not be related to the previous news item.
The shifts are on the order of a pixel, but appear to be elastic. In other
words comparisons taken before and after slews would not necessarily show any
shifts, but those taken during a slew may show shifts. Observers should refrain
from taking calibrations during telescope slews.
Some evidence of wavelength instability. (Mar 5 2006)
Normally, once grating angles are set to select a particular
wavelength region, HIRES is extremely stable.
Recently though, we have received reports
that drifts in the echelle direction, on the order of 0.1 pixel per hour have
been measured. Occasionally, these drifts are monotonic over the course of a
night resulting in drifts on the order of a full pixel. An investigation is
pending. In the mean time, observers may wish to take calibrations
occasionally during the night.
Newly determined throughput curves. (Dec 22 2005)
Throughput curves for the new detector with both cross dispersers
have recently been measured. The curves and a description of how they were
obtained can be found here.
Newly determined values for gain and read noise. (Dec 5 2005)
The gains and read noise in low gain mode have recently been
redetermined. The gains are 1.95, 2.09, and 2.09 electrons/ADU for the blue,
green, and red chips (CCD1,CCD2,CCD3) respectively. The corresponding
read noise values are 2.9, 3.0, and 3.1 electrons.
QE variations on middle CCD (update). (Sep 21 2005)
Work to recover the QE on middle CCD has finished and preliminary
indications are that the QE of all three chips is uniform and back to previous
values. There have been a small number of changes to various quantities such as bias levels and best spectrograph focus though.
Further information, can be found
QE variations on middle CCD. (1 Aug 2005)
We have become aware recently of variations in the QE across the
middle CCD of the HIRES 3-chip mosaic. It appears at this writing that the
variations depend on wavelength, have arisen within the past couple months,
have amplitudes of up to 20 percent. The variations appear to flat field out.
Further information, including plans to resolve this problem can be found
New version of user notes. (18 May 2005)
Observers are reminded that a user notes document is available
via the Observing section of the HIRES pages. As we learn more
about the new detector this is the best place to find information
until a new manual is written.
Changes to quartz lamp(s). (12 April 2005)
Until now, only a single quartz flat field lamp was available to
observers. The drop down lamp selection menu on XHIRES now allows observers to
choose either quartz1 or quartz2.
In the event one burns out during the night there is now a hot spare.
This change means that scripts which previously turned on a lamp simply called
quartz will need to be changed.
Web version of echelle simulator. (5 Feb 2005)
Rather then expend limited resources to fix the bugs described
below in the new echelle simulator, we are working on developing a new version
that is available via the web. A preliminary version is available
Changes to default decker plates. (10 Jan 2005)
As noted below, a new decker plate with 0.4 arcsec slits is
available. This plate replaces the "A" plate which was intended for use
with the adjustable slits. The "A" plate is now only available under
exceptional circumstances. Those considering its use should contact their
Bugs in new echelle simulator. (13 Sep 2004)
We are working to adapt the HIRES echelle simulator to the new dete
ctor, but two shortcomings exist still. There is a sign flip in the echelle an
gle delivered by the simulator. The reported x-coordinate also has a flip in t
he direction it indicates is positive. Please contact your support astronomer
for further details.
New detector improves resolution. (7 Sep 2004)
Smaller pixels and improved CCD characteristics have led to better
resolution. In this plot several Th/Ar lines are co-
added to give an indication of the improvement. For the 0.574 arcsec decker pl
ate shown, the implication is that the new detector improves resolution by about
10 or 15 percent. New decker plates with 0.4 arcsec slits are available which
offer resolution in the 80,000 to 90,000 range.
HIRES PIs can now access data through the Keck Observatory Archiv
e (2 Sept 2004)
Relative QE curve for new detector available (29 Aug 2004)
Observations of the spectrophotometric standard star BD+28d4211 usi
ng identical spectrograph setups have allowed us to compare the relative respons
e of the new mosaic detector to the old detector. In this plot the points are colour coded according to which of the three new ch
ips the slitless spectra were measured on.
Keck Observatory Archive (KOA) sees first bits! (25 Aug 2004)
With the ingestion of the data obtained with the new HIRES detector
during commissioning, the Keck Observatory Archive (KOA) has begun
its operations. PIs will soon receive instructions on how to access
their proprietary data via the KOA user interface. To facilitate the
proper archiving of data, please be sure to use the HIRES number
account provided at each observing run.
Upgraded HIRES detector sees first light! (25 Aug 2004)
The long awaited new HIRES detector saw first light on August 18
and preliminary results indicate it is performing as hoped for.
Improvements in QE, resolution, CTE, spectral coverage, readout
noise etc. are all evident. Please watch this space for further
New HIRES calibration tool available (23 May 2004)
Observers are now provided with a GUI for taking calibrations.
This GUI allows all flats, biases, darks, and comparisons to be
taken in one go. The tool will also shut down the instrument
after the exposures finish if desired.
Th/Ar lamp number 1 should not be used (10 March 2004)
ThAr1 lamp is fading (failing?) again. Observers are
cautioned not to select this lamp (ThAr1), when choosing a
comparison lamp. The other Th/Ar lamp (ThAr2) appears
to be normal in all respects, and should be used.